OpenStudy (anonymous):

find the point on the graph of f(x)=x^2 that is closest to (2 1/2)

OpenStudy (amistre64):

2 1/2 ?

OpenStudy (anonymous):

2,1/2

OpenStudy (anonymous):

@amistre64?

OpenStudy (amistre64):

define -1/y': i get -1/(2x) for any given point (a,b), the slope of the line is -1/(2a) you want this to define the line: y = -1/(2a) (x-2)+1/2 or theres a different approach using a circle centered at (2,1.2)

OpenStudy (amistre64):

(x-2)^2 + (y-1/2)^2 - r^2 = x^2 - y

OpenStudy (amistre64):

\[(x-2)^2 + (y-1/2)^2 - r^2 = x^2 - y\] \[x^2-4x+4 + y^2-y - r^2 = x^2 - y\] \[-4x+4 + y^2 - r^2 = 0\] \[r^2=y^2-4x+4\]

OpenStudy (anonymous):

ok lets focus on the calculus based one

OpenStudy (anonymous):

which is the first one you showed me right?

OpenStudy (amistre64):

yes

OpenStudy (anonymous):

ok so I understand the first step, your simply finding the derivative

OpenStudy (anonymous):

The next step I get as well which is the slope

OpenStudy (amistre64):

finding the derivate, and peroping the slope yes f(x)=x^2 f'(x) = 2x ; -1/f' = -1/(2x)

OpenStudy (anonymous):

but then you lost me at the third step though

OpenStudy (amistre64):